# Two-Stage Learning of Stabilizing Neural Controllers via Zubov Sampling and Iterative Domain Expansion

### v1

@@ 3.1.2 CEGIS based Refinement
Many previous works chang2019neural ; wu2023neural ; yang2024lyapunov start their training with a counterexample-guided approach (CEGIS (abate2018counterexample, )). However, without a good initialization, the counterexamples generated cannot be easily eliminated and will accumulate through the training causing failure, which is shown in the ablation studies table 2. We found that our first stage can yield a great initialization for CEGIS-based training. The controller after the pretraining stage is already almost working, meaning that almost all trajectories starting within $V_{\theta}^{\leq c}$ will successfully converge to the equilibrium. However, as we have discussed, no certification can be given at this time since random samples typically are not sufficient to guarantee the condition holds over the entire continuous domain. Therefore, another stage of CEGIS-based finetuning is performed to ensure the certifiability of the system. To find counterexamples, we follow the previous works as in yang2024lyapunov ; wu2023neural to use a cheap gradient-based PGD attack with the objective
@@
Maximizing this objective leads to potential $x_{\text{cex}}$ that violates the Lyapunov condition in the band $V_{\theta}^{\leq c}\setminus V_{\theta}^{\leq c^{\prime}}$. Here we require a lower $c^{\prime}$ since there is no guarantee that $V_{\theta}(0)=0$, and so near the origin there must exist some small violations. However, we can choose this $c^{\prime}$ to be just slightly larger than $V_{\theta}(0)$. We then eliminate these counterexamples with the loss function
$L_{\text{cegis}}(\theta)$ $=\text{ReLU}(\nabla V_{\theta}(x_{\text{cex}})f(x_{\text{cex}}))$ (15)
where it enforces the Lyapunov condition. Moreover, the overall shaping of the Lyapunov function should mostly be formed at the ROA estimation stage due to the presence of Zubov-inspired loss and should thus be only slightly refined at the CEGIS stage. Therefore, we apply the regularization term
@@
with the aim of preventing the levelset of $V_{\theta}$ from changing drastically during CEGIS. The first term stabilizes the levelset structure near the origin by restricting $V_{\theta}(0)$ to stay close to the original value. The second term stabilizes the levelset near the boundary of the training domain and encourages the ROA to stay away from this boundary to prevent trajectories from escaping from it.
@@
1:Network Parameters $\theta$, Starting Domain $\Omega$, Maximum Iterations $M_{1},M_{2}$, Integration Time $T$ and Discretizations Timesteps $dt$, Learning rate $\alpha$, Number of Training Points $N$, Domain Update Frequency $\gamma$, Number of Trajectories $N_{T}$, Lyapunov Upper Threshold $c$, Domain Expansion Factor $\beta$. PGD steps $P$ and step size $\alpha_{1}$, Finetuning learning rate $\alpha_{2}$ and epochs $K$
@@
4:         $\Omega$leftarrow UpdateDomain($\theta$, $\Omega$, $T$, $dt$, $N_{T}$, $c$, $\beta$) $\triangleright$ See algorithm 2
5:     $x\leftarrow$ $N$ data points from (11) and (12)
@@
7:     $\theta\leftarrow\theta-\alpha\nabla L(\theta,x,x_{\text{boundary}},T,dt)$triangleright See appendix section A.3
@@
[…]
@@
13:         $x\leftarrow x+\alpha_{1}\nabla L_{\text{cex}}(x)$triangleright See equation (14)
14:     Dataset $\leftarrow\{\text{Dataset},x\}$triangleright Buffer has length limit. See appendix A.2
@@
16:         $\theta\leftarrow\theta-\alpha_{2}\nabla(\beta_{1}L_{\text{cegis}}(\theta)+\beta_{2}L_{\text{reg}}(\theta))$triangleright See equation (15) and equation (16)
@@ 3.2 Verification
Criterion: Since there is no guarantee that $V_{\theta}(0)=0$, counterexamples near origin are unavoidable. As the Lyapunov condition enforces the Lyapunov function value to decrease along trajectories, we propose to exclude an unverifiable region defined as a sublevel set of the Lyapunov function. If the excluded set is small enough, this verification suffices for most practical needs. Formally, we propose the verification criterion
@@
[…]
@@
The proof of this theorem can be found in appendix section B.1. Here equation (17) states the normal Lyapunov condition has to hold in the band $\{x:c_{1}\leq V(x)\leq c_{2}\}$ whereas equation (18) states that the trajectory cannot escape through the boundary. Compared to previous works that simply exclude a rectangle near the origin yang2024lyapunov , our verification scheme is more principled since it also provides some guarantees inside the unverifiable region. In systems where formal certification over the full domain is challenging, verifying the condition within a thin band where $|c_{2}-c_{1}|<\epsilon$ for arbitrarily small $\epsilon$ still allows one to certify the forward invariance of the sublevel set $V^{\leq c_{1}}$.
@@
Formal Verification: For formal verification, instead of the commonly used SMT solver dReal in previous works (gao2013dreal, ; chang2019neural, ; liu2025physics, ), we hope to use a state-of-the-art neural network verification framework, such as $\alpha,\!\beta$-CROWN (zhang2018efficient, ; salman2019convex, ; xu2020automatic, ; xu2021fast, ; wang2021beta, ; zhang2022general, ), to rigorously verify the Lyapunov condition through the tool’s efficient linear bound propagation and branch-and-bound procedures. However, several key challenges remain for the application of $\alpha,\!\beta$-CROWN. First, since the continuous-time Lyapunov condition involves Jacobian-vector products, it is essential to support efficient and tight bound propagation for the Jacobian of various dynamical system operators, which were previously not supported in any NN verifiers. Second, even after applying a CEGIS procedure, small counterexamples can still arise near the levelset boundaries at $c_{1}$ and $c_{2}$. To identify the tightest verifiable values of $c_{1}$ and $c_{2}$, a length bisection is typically required, as commonly done in prior work (yang2024lyapunov, ; liu2025physics, ; wang2024actor, ).
@@
In our work, we extend $\alpha,\!\beta$-CROWN with the capability of efficiently and tightly bound the Jacobian of many commonly used operators such as Tanh, Sigmoid. We implemented new linear relaxations that are much tighter compared to naively treating their Jacobian as composite functions of already supported operators, where the details can be found in the appendix section B.2. Furthermore, we enable our verifier to dynamically update $c_{1},c_{2}$ during verification to avoid expensive bisection while remain sound. Intuitively, during the branch-and-bound process, we pick out those domains currently with the worst bound and try to find counterexamples in those domains. If counterexamples are found, we update $c_{1}$ and $c_{2}$ to exclude the counterexamples. Notice that after this update, all the verified subproblems during branch and bound still remain sound and do not need to be verified again. This effectively saved the time needed for the repetitive verification needed if $c_{1}$ and $c_{2}$ are to be found by bisection. The full verification algorithm can be found in appendix section B.3.
@@
Verification Schemes: Compared to discrete-time systems, where the final verification is performed on the discrete Lyapunov condition $\Delta V(x)=V(g(x,u(x))-V(x)<0$ through domain. Continuous-time systems require formally verifying the Jacobian vector product (17). These many multiplications are a big challenge for current verifiers since it is hard to bound multiplication well with a linear relaxation. Therefore, even with the strongest verification tool, verification is hard to scale to higher dimensions due to bound tightness issues. We thus define three verification schemes we considered. As system dimensionality increases, we must rely increasingly more on empirical checks.
@@
    Empirical PGD Verification: In this scheme, we used a strong PGD attack to attempt finding counterexamples of the condition in theorem (3.1). If counterexamples are not found, we say the empirical PGD verification succeeds.
@@
    Monte-Carlo Trajectory Verification: In this scheme, we randomly sample trajectories from the sublevel set $V^{\leq c_{2}}$ and determine whether the trajectories converge to the desired equilibrium.
We can notice that these schemes are more and more empirical but are also more and more scalable to higher dimensional systems. The formal verification scheme, even though it is the strongest verification and can imply the other two schemes, can be hard to scale up. The empirical PGD verification can scale up better, but still fails on our hardest 12-dimensional system. However, the Monte-Carlo trajectory verification can be easily scaled up to any dimension. While this sampling approach provides no formal guarantees, several previous works utilize this technique to validate the learned controller (zeng2024data, ; dawson2022safe, ; yang2023approximate, ).
@@ 4 Experiments
4.1 Setups
We demonstrate the effectiveness of our methods in several challenging benchmarks including four 2D systems (Van-Der-Pol, Double-Integrator, Inverted Pendulum, Path Tracking), one 4D system (Cartpole), three 6D systems (PVTOL, 2D Quadrotor, Ducted Fan), and a 12D system (3D Quadrotor). System details follow prior works (wu2023neural, ; yang2024lyapunov, ; shi2024certified, ) and are also provided in the appendix C.2. We compare our training framework against baselines from DITL (wu2023neural, ), Yang et al. (yang2024lyapunov, ), and Shi et al. (shi2024certified, ). We note that all of them work with discrete-time systems. We focus on comparing to these baselines due to limitations in all the existing continuous-time methods. Both (chang2019neural, ) and (wang2024actor, ) incorrectly handle the condition $u(x^{*})=u^{*}$, causing the origin to fail to be the equilibrium, and the convergence guarantee is thus unsound. Further discussion and examples are included in the appendix C.1. For each method, we run the training/verification pipeline five times with different seeds. We compare training success rates and ROA volumes, where the latter is estimated following previous works (wu2023neural, ; shi2024certified, ) as the fraction of points within the ROA sublevel set among a dense sample from the full domain. All models undergo the three verification steps from section 3.2 in reverse order. More experimental details can be found in appendix C.
@@ 4.1 Setups
Figure 2: Region-of-Attraction (ROA) for Double Integrator and Cartpole compared to baselines. The dotted box shows the baseline’s training box being used. For the double integrator system, it is clear that the initial training box is limiting the baseline’s potential but is not constraining us.
4.2 Main Results
Two and Four Dimensional Systems: On all the 2D systems, our methods robustly produce much larger ROAs compared to existing baselines. On both the path tracking and the inverted pendulum system, our training framework yields $\mathbf{5}$ times bigger ROA compared to the baselines as shown in table 1. Moreover, on the more challenging systems such as inverted pendulum with small torque, our method is the only one that can get perfect success rate. Visualizations of the learned ROA for Double Integrator can be seen in the left of figure 2. We also test our method on a more challenging Cartpole system with four dimensions. Compared to the baselines, we obtain an ROA with more than $\mathbf{330\times}$ larger than the previous state of the art as can be seen in table 1. The visualizations for different projections of the ROA for Cartpole can be found on the right of figure 2. The rest visualizations and details can be found in the appendix C.6.
@@ 4.2 Main Results
Table 1: Success rates (Succ) and ROA volume (ROA) across methods. Dashes () denote systems not covered by the baseline and it is not easy to incorporate the system in their methods. The systems with indicates that only the empirical PGD verification is performed, and the system with ∗∗ implies only trajectory based verification was confirmed. The rest are formally verified. "B" in system implies the Big-torque setting, and "S" implies Small-torque.
System DITL (wu2023neural, ) Yang et al. (yang2024lyapunov, ) Shi et al. (shi2024certified, ) Ours
ROA Succ ROA Succ ROA Succ ROA Succ
Van-der-Pol $1.36\pm 1.57$ $80\%$ $20.2\pm 18.3$ $80\%$ $\mathbf{57.6\pm 3.4}$ $100\%$
Double Int $18.18\pm 16.19$ $60\%$ $130.3\pm 3.9$ $60\%$ $\mathbf{302.5\pm 10.7}$ $100\%$
Pendulum B $61\pm 31$ $100\%$ $70.6\pm 12.2$ $100\%$ $487.5\pm 58.5$ $80\%$ $\mathbf{2946.5\pm 149.1}$ $100\%$
Pendulum S $217.34\pm 6.07$ $60\%$ $306.3\pm 48.7$ $40\%$ $\mathbf{1169.2\pm 124.5}$ $100\%$
Path Tracking B $9\pm 3.5$ $100\%$ $24.06\pm 0.29$ $100\%$ $15.3\pm 8.9$ $60\%$ $\mathbf{122.0\pm 3.7}$ $100\%$
Path Tracking S $14.86\pm 0.18$ $100\%$ $12.5\pm 6.5$ $100\%$ $\mathbf{73.8\pm 12.5}$ $100\%$
Cartpole $0.021\pm 0.012$ $100\%$ $0.9266$ $20\%$ $\mathbf{306.1\pm 54.2}$ $100\%$
PVTOL $49.87\pm 3.91$ $80\%$ $\mathbf{(1.91\pm 0.23)\cdot 10^{4}}$ $100\%^{*}$
2D Quadrotor $2.33\pm 0.47$ $100\%$ $44.53\pm 18.38$ $80\%$ $\mathbf{(6.64\pm 4.67)\cdot 10^{6}}$ $100\%^{*}$
Ducted Fan $\mathbf{(4.31\pm 1.86)\cdot 10^{4}}$ $100\%^{*}$
3D Quadrotor∗∗ $\mathbf{(1.17\pm 0.64)\cdot 10^{9}}$ $100\%^{**}$
Verification Time: As we have discussed, our verifier struggles to verify higher-dimensional systems due to the bound tightness issue over Jacobian vector product. However, we shall demonstrate that our approach is much faster compared to the previous state-of-the-art verifier dReal gao2013dreal used by many previous work chang2019neural ; liu2025physics in continuous time systems. In the 2D systems verification, our method yields 4010000×40-10000$\times$40 - 10000 × faster verification compared to dReal. And for the cartpole system, the dReal fails to return a status. The full verification time comparison can be found in table 3.
4.3 Ablation Studies
We also investigate the effectiveness of our training design choices. We mainly focus on the effectiveness of our training sample selection scheme and the two-stage training by considering three training settings:
@@ 4.3 Ablation Studies
We then investigate the success rate of these three training schemes under different verification schemes with two higher dimensional systems Cartpole and 2D quadrotor. The overall result can be found in table 2. We can found that without our careful design, neither training Scheme 1 or Scheme 3 can yield any verifiable result. However, with our designed ROA estimation stage, even without CEGIS our trained controller and Lyapunov function can robustly pass the trajectory-based verification.
Table 2: Success rates of different training schemes under three verification regimes. “Exp.” denotes dynamic training-domain expansion.
System Training scheme Verification scheme (success)
@@
Cartpole Two-stage with Rand. sampling 0 0 0
ROA-Est. stage only (Zubov + Exp.) 0 0 100%
CEGIS only (with Exp.) 0 0 0
@@
2-D Quadrotor Two-stage with Rand. sampling 0 0 0
ROA-Est. stage only (Zubov + Exp.) 0 0 100%
CEGIS only (with Exp.) 0 0 0

### v2

@@ 3.1.2 CEGIS-based Refinement
Many previous works chang2019neural ; wu2023neural ; yang2024lyapunov start their training with a counterexample-guided approach (CEGIS (abate2018counterexample, )). However, without a good initialization, the counterexamples generated cannot be easily eliminated and will accumulate through the training and cause failure eventually, which is shown in the ablation studies Table 2. We found that our first stage can yield a great initialization for CEGIS-based training. The controller after the pretraining stage is already almost working, meaning that almost all trajectories starting within $V_{\theta}^{\leq c}$ will successfully converge to the equilibrium. However, as we have discussed, no certification can be given at this time since random samples typically are not sufficient to guarantee the condition holds over the entire continuous domain. Therefore, another stage of CEGIS-based finetuning is performed to ensure the certifiability of the system. To find counterexamples, we follow the previous works as in yang2024lyapunov ; wu2023neural to use a cheap gradient-based PGD attack with the objective
@@
Maximizing this objective leads to potential $x_{\text{cex}}$ that violates the Lyapunov condition in the band $V_{\theta}^{\leq c}\setminus V_{\theta}^{\leq c^{\prime}}$. Here we require a lower $c^{\prime}$ since there is no guarantee that $V_{\theta}(0)=0$, and so near the origin there must exist some small violations. However, we can choose this $c^{\prime}$ to be just slightly larger than $V_{\theta}(0)$. These counterexamples then serve as the dataset for the CEGIS training. We then eliminate these counterexamples with the loss function
$L_{\text{cegis}}(\theta;x_{\text{cex}})$ $=\text{ReLU}(\nabla V_{\theta}(x_{\text{cex}})f(x_{\text{cex}}))$ (17)
where it enforces the Lyapunov condition. Moreover, the overall shaping of the Lyapunov function should mostly be formed at the ROA estimation stage due to the presence of Zubov-inspired loss and should thus be only slightly refined at the CEGIS stage. Therefore, we denote $v_{0}$ as the value of $V_{\theta}(0)$ prior to the CEGIS finetuning, and apply the regularization term
@@
with the aim of preventing the levelset of $V_{\theta}$ from changing drastically during CEGIS. The first term thus stabilizes the levelset structure near the origin by restricting $V_{\theta}(0)$ to stay close to the original value, whereas the second term stabilizes the levelset near the boundary of the training domain and encourages the ROA to stay away from this boundary to prevent trajectories from escaping from it.
@@
1:Network Parameters $\theta$, Starting Domain $\Omega$, Maximum Iterations $M_{1},M_{2}$, Integration Time $T$ and Discretizations Timesteps dtdt, Learning rate $\alpha$, Number of Training Points NN, Domain Update Frequency $\gamma$, Number of Trajectories $N_{T}$, Lyapunov Upper Threshold $c$, Domain Expansion Factor $\beta$. PGD steps $P$ and step size $\alpha_{1}$, Finetuning learning rate $\alpha_{2}$ and epochs $K$
@@
4:   $\Omega$ $\leftarrow$ UpdateDomain($\theta$, $\Omega$, $T$, dtdt, $N_{T}$, $c$, $\beta$) $\triangleright$ See algorithm 2
5:  $x\leftarrow$ NN data points from (8) and (9)
@@
7:  $\theta\leftarrow\theta-\alpha\nabla L(\theta,x,x_{\text{boundary}},T,dt)$ $\triangleright$ See loss function section.
@@
9:Dataset $\leftarrow\emptyset$ $\triangleright$ Stage 2: CEGIS Refinement
@@
13:   $x\leftarrow x+\alpha_{1}\nabla L_{\text{cex}}(x)$ $\triangleright$ See equation (16)
14:  Dataset $\leftarrow\{\text{Dataset},x\}$ $\triangleright$ Buffer has length limit. See appendix .
@@
16:   $\theta\leftarrow\theta-\alpha_{2}\nabla(\beta_{1}L_{\text{cegis}}(\theta;\text{Dataset})+\beta_{2}L_{\text{reg}}(\theta))$ $\triangleright$ See section 3.1.2
@@ 3.2 Verification
Criterion: Since there is no guarantee that $V_{\theta}(0)=0$, counterexamples near the origin are unavoidable. As the Lyapunov condition enforces the Lyapunov function value to decrease along trajectories, we propose to exclude an unverifiable region defined as a sublevel set of the Lyapunov function. If the excluded set is small enough, this verification suffices for most practical needs. Formally, we propose the verification criterion
@@
$x\in(V^{\leq c_{2}}\setminus V^{\leq c_{1}})\cap\Omega\;\longrightarrow\;\nabla V(x)\,\cdot\,g\bigl(x,u(x)\bigr)\;<\;0,$ (19)
$x\in\partial\Omega\cap V^{\leq c_{2}}\;\longrightarrow\;g\bigl(x,u(x)\bigr)\,\cdot\,\vec{n}(x)\;<\;0,$ (20)
@@
The proof of this theorem can be found in appendix section B.1. Here equation (19) states the normal Lyapunov condition has to hold in the band $\{x:c_{1}\leq V(x)\leq c_{2}\}$ whereas equation (20) states that the trajectory cannot escape through the boundary. Compared to previous works that simply exclude a rectangle near the origin yang2024lyapunov , our verification scheme is more principled since it also provides some guarantees inside the unverifiable region. In systems where formal certification over the full domain is challenging, verifying the condition within a thin band where $c_{2}-c_{1}<\epsilon$ for arbitrarily small $\epsilon$ still allows one to certify the forward invariance of the sublevel set $V^{\leq c_{1}}$.
@@
Formal Verification: For formal verification, instead of the commonly used SMT solver dReal in previous works (gao2013dreal, ; chang2019neural, ; liu2025physics, ), we hope to use a state-of-the-art neural network verification framework, such as $\alpha,\!\beta$-CROWN (zhang2018efficient, ; salman2019convex, ; xu2020automatic, ; xu2021fast, ; wang2021beta, ; zhang2022general, ; zhou2025clipandverify, ), to rigorously verify the Lyapunov condition through the tool’s efficient linear bound propagation and branch-and-bound procedures. However, several key challenges remain for the application of $\alpha,\!\beta$-CROWN. First, since the continuous-time Lyapunov condition involves Jacobian-vector products, it is essential to support efficient and tight bound propagation for the Jacobian of various dynamical system operators, which were previously not supported in any NN verifiers. Second, even after applying a CEGIS procedure, small counterexamples can still arise near the levelset boundaries at $c_{1}$ and $c_{2}$. To identify the tightest verifiable values of $c_{1}$ and $c_{2}$, a lengthy and expensive bisection is typically required, as commonly done in prior work (yang2024lyapunov, ; liu2025physics, ).
@@
In our work, we extend $\alpha,\!\beta$-CROWN with the capability to efficiently and tightly bound the Jacobian of many commonly used operators such as Tanh, Sigmoid. We implemented new linear relaxations that are much tighter compared to naively treating their Jacobian as composite functions of already supported operators, where the details can be found in the appendix section B.3. Furthermore, we enable our verifier to dynamically update $c_{1},c_{2}$ during verification to avoid expensive bisection while remain sound. Intuitively, during the branch-and-bound process, we pick out those domains currently with the worst bound and try to find counterexamples in those domains. If counterexamples are found, we update $c_{1}$ and $c_{2}$ to exclude the counterexamples. Notice that after this update, all the verified subproblems during branch and bound still remain sound and do not need to be verified again. This effectively saves the time needed for the repetitive verification needed if $c_{1}$ and $c_{2}$ are to be found by bisection. The full verification algorithm can be found in appendix section B.4.
@@
Evaluation Schemes: Compared to discrete-time systems, where the final verification is performed on the discrete Lyapunov condition $\Delta V(x)=V(g(x,u(x))-V(x)<0$ through domain. Continuous-time systems require formally verifying the Jacobian vector product (19). The large number of multiplications is challenging for current verifiers since it is hard to bound multiplication well with a linear relaxation. Therefore, even with the strongest verification tool, verification is hard to scale to higher dimensions due to bound tightness issues. We thus define three evaluation schemes we considered. As system dimensionality increases, we must rely increasingly more on empirical checks.
@@
    PGD Evaluation: In this scheme, we use a strong PGD attack to attempt finding counterexamples of the condition in theorem (3.1). If counterexamples are not found, we say the empirical PGD evaluation succeeds.
@@
    Trajectory Evaluation: In this scheme, we randomly sample trajectories from the sublevel set $V^{\leq c_{2}}$ and determine whether the trajectories converge to the desired equilibrium.
We can notice that these schemes are more and more empirical but are also more and more scalable to higher dimensional systems. The formal verification scheme, even though it is the strongest verification and can imply the other two schemes, can be hard to scale up. The empirical PGD evaluation can scale up better, but still fails on our hardest 12-dimensional system. However, the trajectory based evaluation can be easily scaled up to any dimension. While this sampling approach provides no formal guarantees, several previous works utilize this technique to validate the learned controller (zeng2024data, ; dawson2022safe, ; yang2023approximate, ).
@@ 4 Experiments
Setups: We demonstrate the effectiveness of our methods in several challenging benchmarks including four 2D systems (Van-Der-Pol, Double-Integrator, Inverted Pendulum, Path Tracking), one 4D system (Cartpole), three 6D systems (PVTOL, 2D Quadrotor, Ducted Fan), and a 12D system (3D Quadrotor). System details follow prior works (wu2023neural, ; yang2024lyapunov, ; shi2024certified, ) and are also provided in appendix C.2. We compare our training framework against baselines from DITL (wu2023neural, ), Yang et al. (yang2024lyapunov, ), and Shi et al. (shi2024certified, ). We note that all of them work with discrete-time systems. We focus on comparing to these baselines due to limitations in the existing continuous-time methods. The works (chang2019neural, ; wang2024actor, ; edwards2024fossil, ) incorrectly handle the condition $u(x^{\ast})=u^{\ast}$, so that the origin fails to be the equilibrium for most systems. Further discussions and examples are included in the appendix C.1. However, to ensure fair comparison, we still run and present the comparison with Fossil 2.0 (edwards2024fossil, ) for those systems that their method handle correctly. For each method, we run the training/verification pipeline five times with different seeds. We compare training success rates and ROA volumes, where the latter is estimated following previous works (wu2023neural, ; shi2024certified, ) as the fraction of points within the ROA sublevel set among a dense sample from the full domain. All models undergo the three evaluation steps from section 3.2 in reverse order. More experimental details can be found in appendix C.
@@
Figure 2: Region‑of‑Attraction (ROA) for Double Integrator and Cartpole compared to baselines. The dotted box shows the baseline’s training box being used. For the double integrator system, it is clear that the initial training box is limiting the baseline’s potential but is not constraining us.
Two and Four Dimensional Systems: On all the 2D systems, our methods robustly produce much larger ROAs compared to existing baselines. On both the path tracking and the inverted pendulum system, our training framework yields $\mathbf{5}$ times bigger ROA compared to the baselines as shown in table 1. Moreover, on the more challenging systems such as inverted pendulum with small torque, our method is the only one that can get perfect success rate. Visualizations of the learned ROA for Double Integrator can be seen in the left of Figure 2. We also test our method on a more challenging Cartpole system with four dimensions. Compared to the baselines, we obtain an ROA with more than $\mathbf{330\times}$ larger than the previous state of the art, as can be seen in Table 1. The visualizations for different projections of the ROA for Cartpole can be found on the right of Figure 2. The rest visualizations and details can be found in the Appendix C.6.
@@
Table 1: Success rates (Succ) and ROA volume (ROA) across methods. Dashes (–) denote systems that are not already covered by the baseline and we were unable to obtain successful runs. The "Scheme" column indicates the evaluation scheme for each system as described in section 3.2: Formal = formal verification; PGD = empirical PGD evaluation; Trajectory = trajectory evaluation. "B" in a system implies the Big-torque setting, and "S" implies Small-torque.
System Eval Scheme Fossil (edwards2024fossil, ) DITL (wu2023neural, ) Yang et al. (yang2024lyapunov, ) Shi et al. (shi2024certified, ) Ours
ROA Succ ROA Succ ROA Succ ROA Succ ROA Succ
Van-der-Pol Formal $3.05\pm 0.58$ $60\%$ $1.36\pm 1.57$ $80\%$ $20.2\pm 18.3$ $80\%$ $\mathbf{57.6\pm 3.4}$ $100\%$
Double Int Formal $258.1\pm 36.3$ $80\%$ $18.18\pm 16.19$ $60\%$ $130.3\pm 3.9$ $60\%$ $\mathbf{302.5\pm 10.7}$ $100\%$
Pendulum B Formal $106.5\pm 23.5$ $80\%$ $61\pm 31$ $100\%$ $70.6\pm 12.2$ $100\%$ $487.5\pm 58.5$ $80\%$ $\mathbf{2946.5\pm 149.1}$ $100\%$
Pendulum S Formal $285.2\pm 53.3$ $80\%$ $217.34\pm 6.07$ $60\%$ $306.3\pm 48.7$ $40\%$ $\mathbf{1169.2\pm 124.5}$ $100\%$
Path Tracking B Formal $9\pm 3.5$ $100\%$ $24.06\pm 0.29$ $100\%$ $15.3\pm 8.9$ $60\%$ $\mathbf{122.0\pm 3.7}$ $100\%$
Path Tracking S Formal $14.86\pm 0.18$ $100\%$ $12.5\pm 6.5$ $100\%$ $\mathbf{73.8\pm 12.5}$ $100\%$
Cartpole Formal $0.021\pm 0.012$ $100\%$ $0.9266$ $20\%$ $\mathbf{306.1\pm 54.2}$ $100\%$
PVTOL PGD $49.87\pm 3.91$ $80\%$ $\mathbf{(1.91\pm 0.23)\cdot 10^{4}}$ $100\%$
2D Quadrotor PGD $2.33\pm 0.47$ $100\%$ $44.53\pm 18.38$ $80\%$ $\mathbf{(6.64\pm 4.67)\cdot 10^{6}}$ $100\%$
Ducted Fan PGD $\mathbf{(4.31\pm 1.86)\cdot 10^{4}}$ $100\%$
3D Quadrotor Trajectory $\mathbf{(1.17\pm 0.64)\cdot 10^{9}}$ $100\%$
Verification Time: As we have discussed, our verifier struggles to verify higher-dimensional systems due to the bound tightness issue over Jacobian vector product. However, we shall note that our approach is already much faster compared to the previous state-of-the-art verifier dReal gao2013dreal used by many previous work chang2019neural ; liu2025physics for continuous time systems. In the 2D systems verification, our method yields $40-10000\times$ faster verification compared to dReal. And for the cartpole system, the dReal fails to terminate. The full verification time comparison can be found in Table 3.
Ablation Studies: We also investigate the effectiveness of our training design choices. We mainly focus on the effectiveness of our training sample selection scheme and the two-stage training by considering three training settings:
@@
We then investigate the success rate of these three training schemes under different evaluations with two higher dimensional systems Cartpole and 2D quadrotor. The overall result can be found in Table 2. We can found that without our careful design choices, neither training Scheme 1 or Scheme 3 can yield any verifiable result. However, with our designed ROA estimation stage, even without CEGIS our trained controller and Lyapunov function can robustly pass the trajectory-based verification. Then together with CEGIS stage, our full training pipeline consistently yields controller/Lyapunov function that can be either formally verified or pass the PGD attack evaluation.
Table 2: Success rates of different training schemes under three verification regimes.
System Training scheme Evaluation scheme (% success)
@@
Cartpole Two-stage with Random sampling 0 0 0
ROA-Estimation stage only (Zubov + Domain Expansion) 0 0 100%
CEGIS only 0 0 0
@@
2-D Quadrotor Two-stage with Random sampling 0 0 0
ROA-Estimation stage only (Zubov + Domain Expansion) 0 0 100%
CEGIS only 0 0 0